OpenStudy (anonymous):

Need help with Trig identity question Cot(x)=1+Cot(x)/1+Tan(x)

OpenStudy (loser66):

Prove?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

L.S=R.S

OpenStudy (anonymous):

I know Cot(x)= cos(x)/Sin(x)

OpenStudy (anonymous):

need help with R.S

OpenStudy (loser66):

from the right hand side, 1/ convert cot = cos/sin tan = sin/ cos then replace, 2/ make them have the same denominator 3/ simplify to get the left hand sid.

OpenStudy (anonymous):

I got (1+Cos(x)/Sin(x))*(1+Cos(x)/Sin(x)) for the R.S

OpenStudy (anonymous):

then when i did foil i got (1+Cos(x)/Sin(x))+(Cos(x)/Sin(x))+(Cos^2(x)/Sin^2(x))

OpenStudy (anonymous):

loser66 u left me :c

OpenStudy (loser66):

now I go from right hand side, ok?

OpenStudy (loser66):